Calculate the extra compound interest received when R18 400,00 invested at 13,2% per annum for 11,5 years,

interest paid annually, is invested at the same rate but with interest paid monthly.


1
Expert's answer
2021-05-30T21:05:01-0400

convert R as a percent to r as a decimal

r = R/100

r = 13.2/100

r = 0.132 rate per year,


"A = P(1 + \frac{r}{n})^{nt}"

A=final amount

P=initial principal balance

r=interest rate

n=number of times interest applied per time period

t=number of time periods elapsed


interest rate paid annually

"A = 18,400(1 +\frac{ 0.132}{1})^{(1)(11.5)}"

"A = 18,400(1 + 0.132)^{(11.5)}"

"A = \$76,568.78"

"interest=amount-principal"

"=76568.78-18400=58168.78"

"interest = \$58,168.78"


interest paid monthly

"A = 18,400(1 +\frac{ 0.132}{12})^{(12)(11.5)}"

"A = 18,400(1 + 0.011)^{(138)}"

"A = \$83,267.84"

interest=amount-principal

"=83,267.84-18400=64867.84"

"interest = \$64,867.84"


extra compounding interest

"64867.84-58168.78=6699.06"

"=\$6699.06"